# The Vector Grounding Problem

### v2

@@ 5 Theories of Representational Content: Causes and History
Existing theories of basic representation vary in the emphasis they place on causal-informational or historical relations, as well as in the specific types of such relations they prioritise. Additionally, they may diverge on the inclusion of further relations, such as structural resemblance relations Shea (2018). Nevertheless, all of these theories, in one way or another, assert that a combination of these relations is central to the emergence of meaning in the world. To possess intrinsic meaning, at a minimum, is to stand in a specific set of causal-informational and historical relations with the thing in the world being represented. Theories of content, therefore, provide the essential ingredients that determine referential grounding, enabling internal states to hook onto things in the world. [^18] Philosophical discussions about cognitive representation have traditionally given considerable weight to metaphysical worries regarding the determinacy of content. Once seen instead as scientific, explanatory theories, we take such worries to be misguided (cf., Shea 2018, Coelho Mollo 2022).
@@ 6 Grounding Generative Pre-Trained Models
Philosophical theories of representation identify sets of relations that imbue representations with content, giving them intrinsic meaning. These primarily include causal-informational and historical relations. Our aim is to show that Generative Pre-Trained Models, such as LLMs, can have internal states and generate outputs with intrinsic meaning, i.e., that are referentially grounded in the world. [^19] For recent discussions that apply theories of representational content to LLMs, see Harding (2023), Grzankowski (2024), Goldstein & Levinstein (2024), Chalmers (2025). To begin, we will examine whether appropriate causal-informational relations exist between LLM representations/outputs and entities and properties in the world.
@@ 6.1 Causal-informational relations
A strict separation of form and meaning has been a central assumption in much of linguistics, from Ferdinand de Saussure’s principle of the “arbitrariness of the linguistic sign” (de Saussure 1916) to contemporary formal theories of syntax and semantics (Chomsky 1957, Montague 1970). However, this notion has faced growing criticism since the 1980s, especially in light of empirical investigation of the iconicity of language (Perniss & Vigliocco 2014). In this context, iconicity refers to a non-arbitrary relationship between the form of a linguistic sign and its referent, based on some similarity or analogy (Haiman 1985). The nature of this similarity relation can vary: imagistic iconicity involves a perceptual resemblance between a word and its referent (e.g., in onomatopoeias), while diagrammatic iconicity involves a structural resemblance .
@@
Diagrammatic iconicity is pervasive in linguistic structures. For example, the ordering of events in narrative sequences tends to reflect closeness and succession in time (e.g. “Veni, vidi, vici”); what is nearest to the speaker in a literal or metaphorical sense tends to be mentioned first (e.g., “here and there”); elements whose referents are closely related in some way or other tend to be adjacent in sentences (principle of adjacency); the use of subordination reflects causality or conditionality between states-of-affairs (e.g., “If you study hard, you will pass the exam”), and so on (Van Langendonck 2010). The pervasiveness of diagrammatic iconicity in linguistic structures suggests that the way in which linguistic signs are organised and structured is not arbitrary, and reflect some ways in which the world is organised and structured.
@@
While diagrammatic iconicity is not typically characterised in distributional terms, patterns of co-occurrence in language reflect stable statistical regularities in the way in which linguistic elements are used together. Certain words tend to occur together more frequently than others, and certain grammatical constructions tend to co-occur with particular semantic categories. These patterns of co-occurrence reflect the underlying structure of the world that speakers are trying to describe. Thus, contra Bender & Koller (2020), form and meaning are far from being so easily separable, especially at the level of general patterns of language use.
@@
The patterns of language use derived from the training corpus reveal important aspects of the world. They embed information about patterns of interaction between humans and the world, human ways of categorising their environments, and the selection of features that are more or less salient to them. Indeed, a plausible hypothesis for LLMs’ success in a wide range of tasks is that they have induced general semantic patterns from their training data, which closely resemble what could be termed semantic ‘laws’. [^20] For some preliminary work on the potential form of such semantic laws, see Wolfram (2023). These laws are not merely formal or syntactic; rather, they tap into the intricate relationships between linguistic form and the world it describes. [^21] See Merrill et al. (2024) for concrete evidence that LLMs can induce entailment relationships between sentences from co-occurrence statistics.
@@
In short, LLMs are trained on data that is shaped by humans and their interactions with the world. LLMs stand in no direct causal connection to worldly entities, but they do stand in indirect ones, mediated by the humans who generated the training data. Importantly, while the complete lack of direct causal relations to the world sets LLMs apart from humans, reliance on indirect causal relations does not. Many human representations are also acquired by indirect, mediated ways. Indeed, testimony and social learning underlie a considerable amount of the internal representations humans come to learn, and are considered to be key processes that help explain cultural knowledge and cultural evolution – in their turn central factors for the species’ success Tomasello (2009), Sterelny (2014), Heyes (2018).
@@
These include representations of entities one has not had any direct causal contact with. For example, most people have never been in direct causal contact with enriched uranium, but that does not entail that their thoughts and statements about uranium lack referential grounding – they are indeed about uranium. Representations of culturally constructed notions, such as ‘democracy’ or ‘university’, are also acquired this way – they rely on complex sets of social practices and conventions that are transmitted not by direct causal contact with documents or buildings, but rather by social learning.
@@
In a sense, thereby, the causal-informational relationships between LLMs and the world are dependent on humans. This dependence is however distinct from the one that drives the Vector Grounding Problem. The Problem is not how the representations of connectionist models come to acquire intrinsic meaning, but whether these representations have intrinsic meaning at all – instead of being merely ascribed meaning by external interpreters. To take once again the example of enriched uranium, LLMs find themselves in a position similar to most humans: their representations and outputs about enriched uranium intrinsically refer to enriched uranium – irrespective of any interpreter’s gloss – even though the process enabling such reference depends on a chain of transmission that involves other systems with intrinsic meaning. In our terminology, such internal states are non-basic representations with intrinsic meaning.
@@
In sum, LLMs possess the necessary elements to fulfil the causal-informational requirement for referential grounding. They stand in complex, mediated causal-informational relations to the world. These connections are mediated since they pass through human linguistic production, where language primarily functions as a tool to express aspects of the world. [^22] This analysis would extend, mutatis mutandis, to multimodal inputs, which are also mediated by human interactions with the world (e.g., photographs are captured by humans with a camera). However, one might argue that some worldly properties instantiated by visual samples in a multimodal dataset, such as shapes, can be accessed more directly than they could through the mediation of language. This is because images (and samples in other sensory domains) are entirely iconic, while language only exhibits limited patterns of iconicity. We discuss whether multimodal models fulfil the historical requirement for representational content in subsection 7.3 below. They are complex due to the intricate causal-informational relationships between the patterns of language use, the semantic ‘laws’ that govern them, the structure of the world they have (culturally) evolved to capture, and the causal chains of communication and testimony that produce the learning data – of which the LLM itself can be considered a link. In light of these considerations, we contend that Generative Pre-Trained Models, such as LLMs, can meet the causal-informational criteria necessary for possessing referentially grounded representations and outputs.
6.2 Historical Relations
Although it is important to differentiate between causal-informational and historical relations when considering referential grounding, these aspects are often intertwined in real-world cases. The causal-informational relations previously discussed with respect to LLMs possess a historical dimension, as they depend on patterns of interaction between humans and their environments. Likewise, in many of the interesting cases pertaining to human cognition, the meaning-endowing causal-informational relationships can be traced back to past episodes of learning. That said, as discussed in section 5, finding causal-informational relations between LLMs and the world is the easiest part of our investigation, given how abundant such relations are. Much more challenging is to show that LLMs have internal states that not only carry information about worldly entities, but that also have the function to carry such information – so that the LLM can use it in generating appropriate outputs.
Selection processes and learning are the historical mechanisms that theories of representation commonly invoke to explain how representations come to have the relevant function. These processes explain how a system acquires a specific function or purpose, in this case, the production and/or use of representations. Such functions clarify how cognitive systems can represent accurately – when they produce and use representations that fulfil the relevant function – as well as misrepresent, when the representations produced and/or used are inappropriate for the task. While it is widely accepted that Generative Pre-Trained Models undergo a learning-like process during training, it is less clear whether they undergo the relevant kind of selection process. It can be argued that they do, at least to the same degree as other human artefacts (Artiga 2016, Coelho Mollo 2019), given that numerous design and implementation choices form an integral part of creating artefacts such as AI systems.
The crux of the matter, however, lies in determining whether LLMs stand in historical relations to the right sorts of things, namely the entities and properties in the world that could ground their representations Butlin (2021). At first glance, the situation appears analogous to that of causal-informational relations: LLMs stand in historical relations to the world, albeit in a mediated manner, through the human-generated data used for training them. The critical question is: can the internal states and outputs of LLMs acquire the function to represent extra-linguistic entities, despite their being trained only on language? In other words, the question is whether LLMs can acquire the sort of world-involving functions needed for referential grounding, and which underlie the descriptive normativity that is an essential feature of meaning.
@@ 6.2 Historical Relations
To address this question, it is important to distinguish between the proximate and ultimate functions performed by LLMs during training and inference. The proximate function of LLMs is plainly next-token prediction: predicting the next token in a text sequence, given the context provided by that sequence. This proximate function is infra-linguistic, and so are the standards for its fulfilment or failure, namely, whether the LLM predicts the correct tokens reliably enough. While such function may be suitable for underpinning relational grounding, it appears inadequate for contributing to the world-involving normativity necessary for referential grounding. In contrast, ultimate functions refer to the broader, more abstract goals that the model is tasked with achieving through next-token prediction. These functions can be world-involving in principle, such as generating text that accurately describes real-world states of affairs, or solving problems based on real-world information. In what follows, we will consider several potential sources for world-involving ultimate functions in the training and use of LLMs.
@@
Our general argument goes as follows. The training process of LLMs is a form of selection – through learning – that stabilises the functional roles of the model’s internal states. This process selects for internal states whose ultimate functional role extends beyond mere linguistic prediction to include extra-linguistic, world-involving functions. This is most evident in models that undergo fine-tuning to satisfy human preferences, since such fine-tuning explicitly involves selecting internal states that increase the probability of outputs satisfying world-involving norms, such as factual accuracy. However, we will also present evidence suggesting that pre-training alone can, in some contexts, select for internal states with world-involving content, albeit in a more indirect way.
If LLMs indeed have internal states with extra-linguistic intrinsic meaning, and if those states causally influence the generation of linguistic outputs, then those outputs are thereby referentially grounded in the world. This would mean that at least some outputs of LLMs are intrinsically meaningful in the sense discussed earlier: their meaning does not depend on the interpretations projected onto them by human users.
The key premise in this argument – that the training process of LLMs can select for internal states with extra-linguistic functional roles – requires detailed defence. In what follows, we offer two complementary arguments supporting this premise. The first, which we call the argument from post-training, focuses on how preference tuning methods explicitly establish world-involving functional roles for LLM internal states. The second, which we call the argument from pre-training, examines how even the basic next-token prediction learning objective might, under certain conditions, implicitly select for ultimate world-involving functions.
@@ 6.2.1 The Argument from Post-Training
The training process of modern LLMs, particularly those deployed in consumer-facing chatbots like ChatGPT or Claude, involves two distinct phases: pre-training and post-training. Pre-training involves training the base model from scratch on a next-token prediction objective across vast corpora of text. By contrast, post-training adapts this base model to behave in ways that go beyond generating merely plausible text completions, making LLMs more useful for various downstream tasks and applications. The post-training phase typically includes two types of fine-tuning: instruction tuning, and preference tuning.
Instruction tuning is a form of supervised fine-tuning on curated instruction-response pairs (Ouyang et al. 2022). Given an instruction, the model is trained to generate the desired output by minimising the difference between its predictions and the target responses in the instruction dataset. This teaches the model to generate appropriate responses that follow user instructions rather than simply generating plausible completions of text sequences.
Preference tuning, by contrast, trains LLMs to generate outputs that better align with selected normative criteria, such as helpfulness, harmlessness, and honesty or factual accuracy (Askell et al. 2021). Unlike standard supervised fine-tuning, which only demonstrates what to do, preference tuning explicitly trains models to distinguish between better and worse responses according to these criteria. This is typically implemented by training on pairs of responses where human evaluators have indicated a preference for one over the other.
@@
Reinforcement Learning from Human Feedback (RLHF) is the most well-known preference tuning method (Christiano et al. 2017). RLHF begins by collecting human preferences over model outputs for a diverse set of prompts. These preferences are used to train a reward model that predicts how humans would evaluate various responses. The LLM is then fine-tuned using reinforcement learning to maximize this reward function while remaining reasonably close to its original behaviour. Another increasingly popular approach is Direct Preference Optimization (DPO), which achieves similar results without explicitly modelling the reward function (Rafailov et al. 2024). Instead, DPO directly optimises the model to assign higher probabilities to preferred responses and lower probabilities to rejected ones.
When an LLM is optimised according to human preferences that incorporate extra-linguistic normative standards like factual accuracy, this creates a selection process that reinforces specific activation patterns based on their relationship to worldly states. These internal states acquire a specific functional role: to track relevant aspects of the world that enable the model to generate outputs that satisfy human preferences for truthfulness.
To illustrate how preference tuning establishes world-involving functions, consider a toy example focused solely on factual accuracy using DPO. We begin with a language model that has undergone both pre-training and instruction tuning. This model can follow instructions reasonably well, but has not been specifically optimised for factual accuracy. To align the model with the norm of factual accuracy, we collect a dataset of human preferences where judges rank responses based solely on which one contains more truthful information and fewer falsehoods.
When we apply DPO to this model, it adjusts the model’s parameters to increase the probability of generating factually accurate responses and decrease the probability of generating less accurate ones. [^23] Note that DPO also incorporates a constraint that prevents the model from deviating too far from its original behaviour, to maintain the right balance between preference satisfaction and fluency. Through this process, the model’s internal states are selected specifically because they enable the model to satisfy the norm of factual accuracy, thereby providing a standard of success that depends not on linguistic patterns alone, but on the correspondence between the generated text and states of affairs in the world.
This functional role provides another piece of the story – in addition to causal-informational relations – of how LLM internal states can be referentially grounded, and thus overcome the Vector Grounding Problem. The internal states of the preference-tuned LLM represent aspects of the world because they were selected specifically for their ability to carry information about those aspects in service of generating truthful outputs. When the model succeeds in generating a factually accurate response, its internal representations are fulfilling their function of tracking relevant worldly states.
Importantly, the accuracy of the model is not what determines whether its internal states are referentially grounded. Even a partially inaccurate model prone to occasional hallucinations can have internal states with world-involving content, provided those states were selected for their ability to carry information about aspects of the world. What matters for grounding is not perfect accuracy, but rather the presence of a representational function with world-involving correctness conditions. Preference tuning thus provides a clear mechanism by which LLMs can acquire internal states with extra-linguistic functional roles, insofar as RLHF or DPO incorporate factual accuracy as an explicit criterion over rewarded preferences. [^24] At least in those domains where such factual accuracy has been emphasized during preference tuning.
@@
While we focus on factual accuracy, all three common norms in preference tuning – the “3 Hs” of helpfulness, harmlessness, and honesty – are arguably world-involving. Helpfulness requires generating outputs that genuinely assist users with their goals, which depends on real-world task success conditions rather than mere linguistic coherence. Harmlessness involves avoiding outputs that could lead to real-world negative consequences, requiring sensitivity to causal relationships between linguistic outputs and potential worldly outcomes. Even instruction tuning alone could potentially establish world-involving functions, as successfully following instructions often requires tracking aspects of the world that would make the instruction completion successful. Many instructions implicitly contain success conditions that depend on extra-linguistic factors. Be that as it may, factual accuracy provides the most direct and straightforward case of world-involving content, where the connection between internal states and worldly conditions is the most evident. It offers the clearest case for how LLM internal states can be referentially grounded, and thus endowed with intrinsic meaning.

### v3

@@ 5 Theories of Representational Content: Causes and History
Existing theories of basic representation vary in the emphasis they place on causal-informational relations and selection processes, as well as in the specific types of such relations they prioritise. Additionally, they may diverge on the inclusion of further relations, such as structural resemblance relations (Shea2018). Nevertheless, all of these theories, in one way or another, assert that a combination of these relations is central to the emergence of meaning in the world. To possess intrinsic meaning, at a minimum, is to stand in a specific set of causal-informational and historical relations (via past instances of selection) with the thing in the world being represented. Theories of content, therefore, provide the essential ingredients that determine referential grounding, enabling internal states to hook onto things in the world.
@@ 6 Grounding Generative Pre-trained Models
Philosophical theories of representation identify sets of relations that imbue representations with content, giving them intrinsic meaning. These primarily include causal-informational relations and histories of selection. Our aim is to show that generative pre-trained models, such as LLMs, can have internal states and generate outputs with intrinsic meaning, i.e., that are referentially grounded in the world. [^20] For recent discussions that apply theories of representational content to LLMs, see hardingOperationalisingRepresentationNatural2023c, grzankowskiRealSparksArtificial2024, goldsteinDoesChatGPTHave2024, chalmersPropositionalInterpretabilityArtificial2025, williams2025structuralcorrespondencesgroundreal. To begin, we will examine whether appropriate causal-informational relations exist between LLM representations/outputs and entities and properties in the world.
@@ 6.1 Causal-informational relations
A strict separation of form and meaning has been a central assumption in much of linguistics, from the principle of the “arbitrariness of the linguistic sign” (Saussure1916) to contemporary formal theories of syntax and semantics (chomskySyntacticStructures1957, montagueUniversalGrammar1970). However, this notion has faced growing criticism since the 1980s, especially in light of empirical investigation of the iconicity of language (pernissBridgeIconicityWorld2014). In this context, iconicity refers to a non-arbitrary relationship between the form of a linguistic sign and its referent, based on some similarity or analogy (haimanIconicitySyntax1985). The nature of this similarity relation can vary: imagistic iconicity involves a perceptual resemblance between a word and its referent (e.g., in onomatopoeias), while diagrammatic iconicity involves a structural resemblance.
@@
Diagrammatic iconicity is pervasive in linguistic structures. For example, the ordering of events in narrative sequences tends to reflect closeness and succession in time (e.g., “Veni, vidi, vici”); what is nearest to the speaker in a literal or metaphorical sense tends to be mentioned first (e.g., “here and there”); elements whose referents are closely related in some way or other tend to be adjacent in sentences (principle of adjacency); the use of subordination reflects causality or conditionality between states of affairs (e.g., “If you study hard, you will pass the exam”), and so on (vanlangendonckIconicity2010). The pervasiveness of diagrammatic iconicity in linguistic structures suggests that the organization and structure of linguistic signs is not arbitrary but rather reflects some ways in which the world itself is organised and structured.
@@
While diagrammatic iconicity is not typically characterised in distributional terms, patterns of co-occurrence in language reflect stable statistical regularities in the way in which linguistic elements are used together. Certain words tend to occur together more frequently than others, and certain grammatical constructions tend to co-occur with particular semantic categories. These patterns of co-occurrence reflect the underlying structure of the world that speakers are trying to describe. Thus, contra Bender2020, form and meaning are far from being so easily separable, especially at the level of general patterns of language use.
@@
The patterns of language use derived from the training corpus reveal important aspects of the world. They embed information about patterns of interaction between humans and the world, human ways of categorising their environments, and the selection of features that are more or less salient to them. Indeed, a plausible hypothesis for LLMs’ success in a wide range of tasks is that they have induced general semantic patterns from their training data. These patterns are not merely formal or syntactic; rather, they tap into the intricate relationships between linguistic form and the world it describes. [^21] See merrillCanYouLearn2024 for concrete evidence that LLMs can induce entailment relationships between sentences from co-occurrence statistics.
@@
In short, LLMs are trained on data that is shaped by humans and their interactions with the world. LLMs stand in no direct causal connection to worldly entities, but they do stand in indirect ones, mediated by the humans who generated the training data. Importantly, while the complete lack of direct causal relations to the world sets LLMs apart from humans, reliance on indirect causal relations does not. Many human representations are also acquired by indirect, mediated ways. Indeed, testimony and social learning underlie a considerable amount of the internal representations humans come to acquire, and are considered to be key processes that help explain cultural knowledge and cultural evolution—in their turn central factors for the species’ success (Tomasello2009, Sterelny2014, Heyes2018).
@@
These include representations of entities one has not had any direct causal contact with. Most people have never been in direct causal contact with enriched uranium, but that does not entail that their thoughts and statements about uranium lack referential grounding—they are indeed about uranium. Representations of culturally constructed notions, such as ‘democracy’ or ‘university’, are also acquired this way—they rely on complex sets of social practices and conventions that are transmitted not by direct causal contact with documents or buildings, but rather by social learning.
@@
In a sense, thereby, the causal-informational relationships between LLMs and the world are dependent on humans. This dependence is however distinct from the one that drives the Vector Grounding Problem. The Problem is not how the representations of connectionist models come to acquire intrinsic meaning, but whether these representations have intrinsic meaning at all—instead of being merely ascribed meaning by external interpreters. LLMs find themselves in a position similar to most humans: their representations and outputs about, say, enriched uranium intrinsically refer to enriched uranium—irrespective of any interpreter’s gloss—even though the process enabling such reference depends on a chain of transmission that involves other systems with intrinsic meaning. In our terminology, such internal states are non-basic representations with intrinsic meaning.
@@
In sum, LLMs possess the necessary elements to fulfil the causal-informational requirement for referential grounding. They stand in complex, mediated causal-informational relations to the world. These connections are mediated since they pass through human linguistic production, where language primarily functions as a tool to express aspects of the world. [^22] This analysis would extend, mutatis mutandis, to multimodal inputs, which are also mediated by human interactions with the world (e.g., photographs are captured by humans with a camera). We discuss whether multimodal models fulfil the historical requirement for representational content in subsection 7.3 below. They are complex due to the intricate causal-informational relationships between the patterns of language use, the structure of the world they have (culturally) evolved to capture, and the causal chains of communication and testimony that produce the learning data—of which the LLM itself can be considered a link. In light of these considerations, we contend that generative pre-trained models, such as LLMs, can meet the causal-informational criteria necessary for possessing referentially grounded representations and outputs.
6.2 Selection history
Finding causal-informational relations between LLMs and the world is the most straightforward part of our investigation, given the abundance of such relations. The far greater challenge lies in showing that LLMs have internal states that not only carry information about worldly entities and properties, but also have the function of carrying such information—enabling LLMs to generate appropriate outputs based on this representational content.
Theories of representation typically invoke selection processes, broadly construed to include learning, as the historical mechanisms that explain how internal states acquire representational functions. While generative pre-trained models clearly undergo a form of learning during training, whether this constitutes the appropriate kind of selection process is less clear. The crucial question is whether LLMs’ internal states possess the right kind of selection history—having been selected for carrying information about something and guiding behaviour accordingly—directed toward the right targets, namely the entities and properties in the world that could ground their representations (Butlin2021). Only when these conditions are satisfied can LLMs’ internal states acquire genuine representational functions and the descriptive normativity that accompany them.
@@ 6.2 Selection history
It is important to distinguish between the proximate and ultimate functions that LLMs perform during training and inference. The proximate function of LLMs is plainly next-token prediction: given a text sequence, predict the subsequent token based on contextual information. This proximate function is intra-linguistic, and so is its success criterion—namely, the LLM’s accuracy at predicting the correct tokens. While this function may support intra-linguistic relational grounding, it appears insufficient for establishing the normativity required for genuine reference to the external world. Ultimate functions, by contrast, encompass the broader, more abstract goals that models achieve through next-token prediction. These functions can, in principle, involve world-directed tasks such as generating text that accurately describes real-world states of affairs or solving problems requiring real-world knowledge. We will examine several potential sources of world-involving, representational ultimate functions in LLM training and deployment. [^23] The relevant world need not be our shared, physical world, but can encompass other extra-linguistic domains, such as the ‘game world’ of chess rules, pieces, and moves. When discussing this more liberal notion of ‘world-involving’, we note it explicitly in the text.
@@
Our general argument goes as follows. The training process of LLMs is a form of selection through learning that picks out and stabilises those internal states that lead to successful behaviour. In many cases, this process selects for internal states whose ultimate function extends beyond mere linguistic prediction to include extra-linguistic, world-involving correctness conditions, such as representing facts about the world. This is most evident in models that undergo fine-tuning to satisfy human preferences, since such fine-tuning explicitly involves selecting internal states that increase the probability of outputs satisfying world-involving norms, such as factual accuracy. We will also investigate the possibility that, to a limited extent and within some constrained domains, pre-training on next-token prediction alone may be sufficient for endowing internal states with world-involving, representational functions. If LLMs indeed have internal states with extra-linguistic intrinsic meaning, and if those states are causally responsible for the generation of linguistic outputs, then those outputs are thereby referentially grounded in the world. This would mean that at least some outputs of LLMs are intrinsically meaningful in the sense discussed earlier: their meaning does not depend on the interpretations projected onto them by human users.
The key premise in this argument—that the training LLMs undergo is a kind of selection process suitable for endowing their internal states with functions to represent states of the world—requires careful defence. In what follows, we explore three complementary strategies to defend the premise, each aimed at showing that a specific kind of learning process in current LLMs can in principle confer world-involving, representational functions to their internal states. The argument from post-training focuses on fine-tuning methods based on human preferences as the most promising candidates for this role. The argument from pre-training investigates whether the basic next-token prediction objective might, under specific conditions, implicitly select for world-involving representational functions. Finally, the argument from in-context learning proposes that LLMs could, in principle, acquire world-involving representational functions during inference without parameter updates.
@@ 6.2.1 The Argument from Post-Training
The training process of modern LLMs, particularly those deployed in consumer-facing chatbots like ChatGPT, involves two distinct phases: pre-training and post-training. Pre-training involves training the base model from scratch on a next-token prediction objective across vast corpora of text. By contrast, post-training adapts this base model to behave in ways that go beyond generating merely plausible text completions, making LLMs more useful for various downstream tasks and applications. This phase typically includes two types of fine-tuning: instruction tuning and preference tuning.
@@
Instruction tuning is a form of supervised fine-tuning on curated instruction-response pairs (ouyangTrainingLanguageModels2022). Given an instruction, the model is trained to generate the desired output by minimising the difference between its predictions and the target responses in the instruction dataset. This teaches the model to generate appropriate responses that follow user instructions rather than simply generating plausible completions of text sequences. Preference tuning further trains LLMs to generate outputs that better align with selected normative criteria, such as helpfulness, harmlessness, and honesty or factual accuracy (askellGeneralLanguageAssistant2021). Unlike standard supervised fine-tuning, which only demonstrates what to do, preference tuning explicitly trains models to distinguish between better and worse responses according to these criteria. This is typically implemented by training on pairs of responses where human evaluators have indicated a preference for one over the other. Reinforcement Learning from Human Feedback (RLHF) is the most well-known preference tuning method (christianoDeepReinforcementLearning2017). RLHF begins by collecting human preferences over model outputs for a diverse set of prompts. These preferences are used to train a reward model that predicts how humans would evaluate various responses. The LLM is then fine-tuned using reinforcement learning to maximise this reward function while remaining reasonably close to its original behaviour. Another increasingly popular approach is Direct Preference optimisation (DPO), which achieves similar results without explicitly modelling the reward function (rafailovDirectPreferenceOptimization2024). Instead, DPO directly optimises the model to assign higher probabilities to preferred responses and lower probabilities to rejected ones.
When an LLM is optimised according to human preferences, a selection process is established that reinforces specific internal states based on their contribution to externally rewarded outputs. Consider a toy example using DPO. We begin with a language model that has undergone both pre-training and instruction tuning. This model can follow instructions reasonably well, but has not been specifically optimised for factual accuracy. To align the model with the norm of factual accuracy, we collect a dataset of human preferences where judges rank responses based solely on which one contains more truthful information and fewer falsehoods. When we apply DPO to this model, it adjusts the model’s parameters to increase the probability of generating factually accurate responses and decrease the probability of generating less accurate ones. Through this process, the model’s internal states are selected because they enable the model to satisfy the norm of factual accuracy, thereby providing a standard of success that depends not on linguistic patterns alone, but on the correspondence between the generated text and states of affairs in the world. This provides the missing piece of the story—in addition to causal-informational relations—of how the internal states of LLMs can be referentially grounded, and thus overcome the Vector Grounding Problem. The internal states of the preference-tuned LLM represent states of the world because they were selected specifically for their ability to carry information about those states in service of generating truthful outputs. When the model succeeds in generating a factually accurate response, its internal states fulfil their function of representing relevant worldly states.
Two clarifications are in order. Firstly, we should distinguish between the function of the learning mechanism and the function of the representational states selected by that mechanism. The designers’ intentions bestow an artefactual function on the learning mechanism (e.g., the preference tuning pipeline has the intended function of updating parameters based on reward signals). However, the function of the model’s internal states is set by the learning process itself. The situation is analogous to artificial selection in biology. A dog breeder’s intention to produce calmer dogs establishes a selection regime, but it is the objective history of differential reproduction that confers the biological function of producing calmness on the selected traits. This function is a fact about the lineage’s history, not the breeder’s mind. Similarly, the designers of an LLM create a selection apparatus, but the autonomous operation of this apparatus on data confers new, selected-effects functions on the model’s internal states. The normativity—the standard for correctness or incorrectness—is supplied by this objective history of selection, not the designers’ intentions.
@@
Secondly, the extra-linguistic content of the model’s representations is determined by the specific worldly conditions that drove their selection, which in the case of preference tuning means whatever properties the reward signal actually and reliably tracks. In that regard, the focus on factual accuracy in our example is a special case. When human evaluators consistently provide higher rewards for factually accurate statements, the selection process picks out and stabilises the model’s internal states that carry relevant factual information about the world, in virtue of their doing so. The other two common norms in preference tuning—helpfulness and harmlessness—are also suitable to selecting referentially grounded internal states. Helpfulness requires generating outputs that genuinely assist users with their goals, which hinges on real-world task success conditions rather than mere linguistic coherence. Harmlessness involves avoiding outputs that could lead to real-world negative consequences, requiring sensitivity to causal relationships between linguistic outputs and potential worldly outcomes. Even instruction tuning alone could establish world-involving functions, as successfully following instructions often requires tracking aspects of the world that would make the instruction completion successful. Many instructions implicitly contain success conditions that depend on extra-linguistic factors.
In summary, post-training establishes the selection processes necessary for endowing LLM internal states with representational functions, which together with causal-informational relations are the key ingredients for intrinsic meaning. It is worth noting that the specific norms used in post-training shape what the internal states represent. Achieving referential grounding does not automatically guarantee that the meanings of the internal representations and outputs of LLMs align with their conventional human meanings. Un-humanlike selection criteria, together with the other un-humanlike features of LLMs, may endow them with ‘deviant meanings’. This is an intriguing possibility that lies beyond the scope of this paper. What matters for our purposes is that LLMs’ internal states and outputs can be intrinsically meaningful, even if such meanings may be alien to humans.